Business 1B - Managerial Accounting - Master Budget Comprehensive Project
Radical Boards, Inc. manufactures and sells a single model The unit cost for ending FG inventory at 12/31/11 was $374.80.
of snowboard, the Vortex. In the summer of 2011, Ignore Work In Process
Iggy Sawdust, CPA, provided data for the 2012 budget:
Direct Material Requirements (per snowboard) The December 2012 Budget Includes the following selected balances:
Wood 5 board feet (b.f.) Cash $      10,000.00
Fiberglass 6 yards Property, Plant, and Equipment $   850,000.00
Direct Labor Requirements 5 hours Current Liabilities $      17,000.00
Long Term Liabilities $   178,000.00
Expected sales during 2012 - 1000 units at $450 per unit Capital Stock $   600,000.00
Expected inventory balances:
1/1/12 12/31/12
Finished Goods 100 200 Retained Earnings at balance at 12/31/11 $   184,980.00
Wood Raw Material 2000 b.f. 1500 b.f.
Fiberglass Raw Material 1000 yards 2000 yards
Required: Prepare the following reports for the year 2012
Other data: Unit Prices 1 Sales Budget in dollars
Wood $30.00 per b.f. 2 Production Budget in units
Fiberglass $5.00 per yard 3 Direct Materials Budget
Direct labor $25.00 per hour 4 Direct Labor Budget
5 Manufacturing Overhead Budget
Variable Overhead $7.00 per dlh 6 Predetermined Overhead Rate
Fixed Overhead $66,000 7 Estimated Overhead to be Applied to One Snowboard
Overhead Driver DL Hours 8 Calculate the cost to manufacture a snowboard
9 Budgeted Ending Inventory for Materials and Finished Goods
Variable marketing costs $250 per sales visits 10 Cost of Goods Sold Budget
Expected sales visits in 2012 30 11 Budgeted Income Statement
Fixed non-overhead costs $30,000 12 Budgeted Balance Sheet
1) I advise doing all of the calculations by hand before you input numbers in the work sheets.
2) Fill in the hatched boxes ONLY.  Some inputs are just numbers.  Some require calculations.  The rest of the cells calculate automatically.
3) Clarification: The balance sheet numbers above are as of the end of year, i.e. all you have to do is determine year-end retained earnings.
Part 8 - Cost to Manufacture One Snowboard Part 10 - Cost of Goods Sold Budget
Direct Material Inputs Beginning Finished Goods Inventory
Wood Add: Direct Materials Used
Fiberglass Direct Labor Added
Direct Labor Inputs Overhead Applied
Overhead Inputs Total Manufacturing Costs $                          -   
   Total Cost of One Snowboard $                                               -    Cost of Goods Available for Sale $                          -   
Less: Ending Finished Goods Inventory $                          -   
Part 9 - Budgeted Ending Inventory for Materials and Finished Goods = Cost of Goods Sold $                          -   
Cost Per
Units Unit Total
Raw Materials
Wood
Fiberglass
Finished Goods     $                                               -   
Total $                                               -   
Answer #1

8.

Cost to Manufacture One Snowboard
Input Cost per input Total
Direct material inputs
Wood 5 $30 $150
Fiberglass 6 5 30
Direct labor inputs 5 25 125
Overhead inputs 5 19 95
Total cost of one snowboard $400

Units produced = 1,000+200-100 = 1,100

Direct labor hours used = 1,100*5 = 5,500 hours

Total overhead cost = ($66,000+5,500*$7) / 5,500 = $19

9.

Budgeted Ending Inventory For Materials and Finished Goods
Unit Cost per unit Total
Raw materials
Wood 1500 $30 $45,000
Fiberglass 2000 5 10,000
Finished goods 200 400 80,000
Total $135,000

10.

Cost of goods sold budget cannot be calculated as data regarding unit cost of beginning raw material inventory not provided.